OpenStudy (amyga):

Can somebody help me? Matthew is 7 years older than twice the age of his sister. If Matthew is 13, how old is his sister? A)3 B)4 C)10 D)33

OpenStudy (maddie82502):

im pretty sure it is d 33

OpenStudy (fedorable):

The correct answer is A

OpenStudy (fedorable):

@maddie82502 @amyga

OpenStudy (amyga):

Thank you.

OpenStudy (fedorable):

medal please?

OpenStudy (fedorable):

thank you

OpenStudy (maddie82502):

ok I get it @Fedorable

OpenStudy (anonymous):

she's 3. you would set it up so it looks like this: 13= 7 + (2x)--now that you have the equation, all you have to do is solve. 13= 7+ (2x) -7= -7 -- in order to isolate the variable, you have to get rid of the constant using the subtraction property of equality, so in this case you'd subtract 7 from both sides. 6= (2x)-- you would be left with 6 equals 2x. you have to again isolate the variable by this time getting rid of the coefficient, you'd o so by doing the inverse operation which is in this case dividing by two on both sides. 6= (2x) /2= /2 3=x

OpenStudy (anonymous):

A

OpenStudy (koalagurl1):

A

OpenStudy (anonymous):

So what is 13 minus 7 and what is half of that? Can I have a medal please? (:

OpenStudy (anonymous):

a